# COBias and Debias: Minimizing Language Model Pairwise Accuracy Bias via Nonlinear Integer Programming

### v1

@@ 6 Discussion
Using Calibration Methods to Mitigate ICL Biases. Previous calibration methods reduce biases caused by prompting. For example, Zhao et al. [4], Zhou et al. [36] use content-free or content-based prompts to measure an offset and calibrate class probabilities by removing it. This may be good for mitigating bias towards a single class, but it does not consider the interaction between class predictions, e.g., pairwise class accuracy bias studied in this work. Another hazard is using prompts to compute the offset, which may cascade biases brought by a certain or several prompt templates, and biases inherited from pretraining may not be reflected. Another line of works does not rely on further prompting. Han et al. [3] fits prototypical clusters on class probabilities and updates predictions by a cluster-label mapping. We share the same spirit of correcting biased predictions without more prompting, but we differ in that our correction weights are explicitly optimized on COBias and accuracy metrics, while the cluster-based decision boundary is implicitly learned.
@@ 7 Conclusion
This work tackles the imbalance in per-class prediction accuracy of language models. We introduce COBias to assess pairwise class accuracy differences, and reveal that LLMs of varied scales and families exhibit large COBias at inference. To minimize COBias and push for higher accuracy, we view debiasing as a combinatorial optimization problem and propose DNIP, which optimizes class correction weights by jointly minimizing evaluation scores by COBias and maximizing accuracy. Empirical evaluations on three LLMs across seven multi-class NLP classification tasks demonstrate that DNIP simultaneously achieves lower test COBias and higher test accuracy over conventional ICL. We advocate that reducing COBias is a direction to gain more accurate, reliable LLM predictions.
Limitations and Future Work. The limitation is that we only evaluate NLP classification tasks. We leave evaluations on open-ended generations and other modalities for future work. Indeed, DNIP is fruitful for debiasing imbalanced class predictions, which can easily extend to vision/multimodal LMs, as long as there is a need for mitigating per-class accuracy differences. This simple inference-time plug-in solution is there for further boosting LM performances.
Appendix A Derivation of SA Algorithm Complexity
@@
Appendix B Additional Heatmaps for Pairwise Accuracy Differences
Figure 7 and 8 demonstrate that more balanced class accuracies are achieved with DNIP for smaller LLMs of Llama-2-7B and GPT-2-XL.
@@ Appendix B Additional Heatmaps for Pairwise Accuracy Differences
Appendix C Additional Ablation Analysis
We show ablation analysis on GPT-2-XL and Llama-2-7B in Figure 9 and 10.
@@ Appendix C Additional Ablation Analysis
Appendix D Full Table with Varying Number of Optimization Examples for DNIP
The full results for all three models across varying optimization set sizes are shown in Table 2 (on the next page).

### v3

@@ 6 Discussion
More Methods That Potentially Improve Accuracy while Reducing COBias. Prior methods tackle the prediction bias problem from different angles. Although COBias is not be explicitly modeled, they are promising to reduce COBias. Therefore, we draw comparisons to two popular categories of debiasing methods, adaptations and calibrations. In particular, with the same class probability vectors as input, we build a single linear adapter (LA) network with adapter dimension of 64, and train it towards the same objective as DNIP; we also follow the batch calibration method [36] to estimate a calibration term using test examples.
Figure 8: Comparisons to more methods.
Figure 8 shows the comparison results. fine-tuning a simple linear adapter does not work well on the class output probabilities. The proposed method outperforms both methods in accuracy and COBias.
@@
Using Calibration Methods to Mitigate ICL Biases. Previous calibration methods reduce biases caused by prompting. For example, Zhao et al. [24], Zhou et al. [36] use content-free or content-based prompts to measure an offset and calibrate class probabilities by removing it. This may be good for mitigating bias towards a single class, but it does not consider the interaction between class predictions, e.g., pairwise class accuracy bias studied in this work. Another hazard is using prompts to compute the offset, which may cascade biases brought by a certain or several prompt templates, and biases inherited from pretraining may not be reflected. Instead, we correcct biased predictions without more prompting, but we differ in that our correction weights are explicitly optimized on COBias and accuracy metrics, while the cluster-based decision boundary is implicitly learned.
@@
Optimization Algorithms for Integer Programming. The classic solvers for integer programming include operational research techniques such as Branch and Bound (BnB), commonly used for linear integer programming problems. However, it is difficult for these methods to solve our DNIP model, which contains discontinuous functions and is non-differentiable. In this case, a series of metaheuristic algorithms are more suitable, solversincluding Simulated Annealing (SA), Genetic Algorithms, Ant Colony Optimization, and Particle Swarm Optimization. These algorithms all belong to the same category, and each meta heuristic has their own pros and cons. Therefore, SA in this paper has a clear purpose of tackling our NIP mathematical model that classical operational research methods cannot solve.
Our Motivation Is Different from Classical Post-hoc Corrections. Some may contend that achieving equitable accuracies across all classes is a well-explored issue in standard machine learning classifiers. However, it is crucial to see that class prediction accuracy imbalances must be understood within their specific context. The accuracy bias in LLMs’ outputs arises from fundamentally different causes than the unequal class accuracies seen in potentially overfitted traditional classifiers. In the former, biases stem from prompts and pre-training, while in the latter, the imbalance is driven by skewness in the supervised training data. This distinction underscores why our approach is uniquely applied to LLMs’ output token class probabilities.
@@ 7 Conclusion
This work tackles the imbalance in class prediction accuracy of language models. We introduce COBias and reveal that LLMs of varied scales and families exhibit large COBias at inference. To minimize COBias and push for higher accuracy, we view debiasing as an inference-time combinatorial optimization problem and propose DNIP, which optimizes class correction weights by jointly minimizing evaluation scores by COBias and maximizing accuracy. Empirical evaluations on three LLMs across seven multi-class NLP classification tasks demonstrate that DNIP simultaneously achieves lower test COBias and higher test accuracy over conventional ICL. We advocate that reducing COBias is a direction to gain more accurate, reliable LLM predictions. For future works, we may extend our studies to more tasks and other modalities. We aim to find simple inference-time plug-in solution for further boosting language model prediction performances.
Appendix A Other Examples of Over-prediction and Under-prediction in LLMs
We present Llama-2-13B’s and ChatGPT’s over- and under-predictions on a class imbalanced classification task, the Drug-Drug Interaction (DDI) relation classification task, where around 90% test instances are Negative. As shown in Table 5, Llama-2-13B significantly under-predicts Negative. In Table 6, similar under-predictions in the DDI Negative class happens for ChatGPT (GPT-3.5-turbo-instruct [39]), the RLHF’d chat LLM with magnitudes of more parameters. (It is noted that at the time of our experiments, ChatGPT does not support returning full probabilities over the entire vocabulary. To investigate the over-/under-prediction issue, we prompt ChatGPT on the DDI task with the following prompt template: “Please choose one most suitable answer for the following description. [ONE-SHOT DEMONSTRATION] Description: [TEST INPUT] Q: What is the drug-drug interaction relation between the aforementioned @DRUG\$ pair? Choose from Negative, Mechanism, Effect, Advice, Interaction. A: ”, and use the answer token directly as the prediction, and plot the confusion matrix in Table 6. For fair comparisons, we prompt Llama-2-13B in the same manner to obtain Table 5, which is different from how we prompt Llama-2-13B in the main paper.)
Pred.
Neg Eff Mech Adv Int
Neg 1749 438 1498 546 506
Eff 118 33 110 52 47
True Mech 77 28 123 45 29
Adv 84 15 62 36 24
Int 34 11 28 13 10
Table 5: Over-prediction and under-prediction in Llama-2-13B, as shown by the confusion matrix on DDI test set.
Pred.
Neg Eff Mech Adv Int
Neg 809 707 1001 217 2003
Eff 45 109 79 6 121
True Mech 15 46 172 0 69
Adv 24 12 18 79 88
Int 9 9 7 1 70
Table 6: Over-prediction and under-prediction in ChatGPT, as shown by the confusion matrix on DDI test set.
The over-/under-predictions in LLMs of different sizes and families, and the inconveniences in training/fine-tuning LLMs, together call for a more actionable bias metric quantifying the over-/under-prediction issue and an accompanying debiasing method tailored for LLMs.
Appendix B Directly Optimizing the Odd Prediction Rate
We discuss another option of measuring class accuracy imbalances, which does not mitigate odd predictions as effective as using COBias, though using it can obtain similar or slightly higher overall accuracy. It is the odd prediction rate, i.e., the rate of odd predictions over the number of ground-truth instances, which is a measurement reflecting the odd class and the level of “accuracy hurt” caused by the odd class. We define $H(c_{i})$ as below.
$\textrm{odd}_{i}=\verb|odd_class|(c_{i}),$ (15)
$H(c_{i})=\frac{\displaystyle\sum\nolimits_{m}\mathbbm{1}\{y_{m}=c_{i},\hat{y}_{m}=\textrm{odd}_{i}\}}{\displaystyle\sum\nolimits_{k}\mathbbm{1}\{y_{k}=c_{i}\}},$ (16)
where $\textrm{odd}_{i}$ is the odd class to $c_{i}$. For example, one may quantify the effect of a current odd class Business to the ground-truth class Tech as the rate of Business predictions over ground-truth Tech instances, i.e., 822/1240, and an immediate debiasing goal is to minimize this rate. The DNIP objective with the odd prediction rate is as follows.
$\min\quad Z(\xi)=\frac{1}{M}\displaystyle\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}+\beta\frac{1}{N}\displaystyle\sum\nolimits_{i}H(c_{i})-\tau\displaystyle\sum_{j=1}^{N}\textrm{PMI}_{j}$ (17)
We compare the effectiveness of using COBias, COBiassingle, and $H$ in the DNIP objective, and refer to the respective objectives as ZCB (Eq. 4), ZCBs (Eq. 14), and ZH (Eq. 17). Detailed test scores for the benchmark datasets on Llama-2-13B, including class accuracies, are shown in Tables 7 to 13.
In short, ZH still leaves some classes under-predicted, as examplified by the individual test class accuracies, and test COBias (or test COBiassingle) scores on TREC, RTE, etc. The rationale is that optimizing bias measurements that extend accuracy differences, such as the odd prediction rates, does not mitigate odd class predictions better than optimizing COBias directly. ZCB most effectively mitigates the over- and under-prediction issue, balancing class accuracies.
Optimizing other extended measurements, such as precision or $F_{1}$ scores differences, may obtain similar performances to the three existing metrics, which will be left for future discussions.
DNIP Objective (Results without DNIP are shown on the last row.) Evaluation Metric
CB CBs H Overall Acc. Class Acc.
Co. Sch. Art. Ath. Polit. Trans. Bldg. Nat. Vill. Anim. Plt. Alb. Film Bk.
ZCB 7.7 7.5 3.3 93.4 82.6 97.7 94.0 99.5 97.9 94.4 86.6 89.0 99.5 96.3 98.5 98.6 97.0 77.8
ZCBs 9.1 5.7 3.9 92.6 87.4 96.8 97.1 98.9 98.1 87.3 87.6 84.0 99.3 99.2 97.9 99.6 96.5 69.0
ZH 8.9 9.5 2.8 93.2 76.2 99.3 95.0 99.4 98.8 96.9 82.8 89.8 99.5 98.4 98.9 99.8 96.1 76.3
ICL 16.2 13.7 5.6 88.6 91.7 99.1 92.7 99.7 97.0 94.8 78.4 44.2 99.7 96.9 99.3 99.8 95.2 63.3
Table 7: Comparison of DBpedia test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
DNIP Objective (Results without DNIP are shown on the last row.) Evaluation Metric
CB CBs H Overall Acc. Class Acc.
Wrd. Sports Bus. Tech.
ZCB 6.3 4.4 7.2 87.9 85.3 95.9 86.5 83.9
ZCBs 7.3 4.3 6.9 88.4 84.8 98.2 85.8 84.7
ZH 8.5 4.8 6.5 88.3 86.8 98.7 84.2 83.2
ICL 28.3 22.5 14.6 79.9 81.5 98.7 91.0 48.0
Table 8: Comparison of AGNews test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
DNIP Objective (Results without DNIP are shown in the last row.) Evaluation Metric
CB CBs H Overall Acc. Class Acc.
Neg. Eff. Mech. Adv. Int.
ZCB 7.5 7.0 39.6 40.4 42.1 31.6 32.8 33.9 31.6
ZCBs 19.9 16.8 42.1 22.9 19.9 34.2 37.3 39.8 44.8
ZH 43.6 44.6 30.9 21.7 17.6 12.6 47.2 85.5 33.3
ICL 45.6 71.8 59.4 7.2 0.0 30.7 61.5 24.1 43.5
Table 9: Comparison of DDI test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
DNIP Objective (Results without DNIP are shown in the last row.) Evaluation Metric
CB CBs H Overall Acc. Class Acc.
Num. Loc. Per. Desc. Ent. Ab.
ZCB 14.2 16.0 18.9 77.1 80.2 92.2 88.7 69.1 64.5 74.1
ZCBs 15.0 13.6 17.7 79.0 90.0 89.3 87.7 70.0 64.5 74.1
ZH 21.3 31.7 12.6 77.3 93.8 88.5 91.3 75.6 39.7 88.9
ICL 35.9 19.9 18.5 68.5 82.9 89.7 88.7 72.9 9.9 96.3
Table 10: Comparison of TREC test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
DNIP Objective (Results without DNIP are shown in the last row.) Evaluation Metric
CB CBs H Overall Acc. Class Acc.
False True
ZCB 4.5 4.5 26.0 74.1 71.8 76.3
ZCBs 4.5 4.5 26.0 74.1 71.8 76.3
ZH 13.3 13.3 24.7 75.7 68.7 82.0
ICL 43.4 43.4 29.7 71.5 48.6 92.0
Table 11: Comparison of RTE test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
DNIP Objective (Results without DNIP are shown in the last row.) Evaluation Metric
CB CBs H Overall Acc. Class Acc.
Trbl. Bad Ok Gd. Grt.
ZCB 18.7 13.2 36.8 48.3 47.0 56.6 19.4 54.1 56.6
ZCBs 29.3 22.5 38.6 49.3 50.7 59.3 0.9 68.9 54.4
ZH 35.6 34.8 34.8 50.6 37.2 70.9 0.9 68.8 53.3
ICL 53.1 63.9 53.1 44.9 28.7 71.6 0.0 81.6 10.8
Table 12: Comparison of SST-5 test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
DNIP Objective (Results without DNIP are shown in the last row.) Evaluation Metric
CB CBs H Overall Acc. Class Acc.
Yes No Mayb.
ZCB 41.1 31.3 38.9 63.1 74.4 61.1 12.7
ZCBs 16.1 17.1 37.1 52.4 56.9 51.5 32.7
ZH 36.3 24.3 27.8 62.7 65.9 72.0 17.6
ICL 61.2 65.0 44.0 55.1 42.2 93.7 1.8
Table 13: Comparison of PubMedQA test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
Appendix C Derivation of SA Algorithm Complexity
@@
Appendix D Additional Heatmaps for Pairwise Accuracy Differences
Figure 9 and 10 demonstrate that more balanced class accuracies are achieved with DNIP for smaller LLMs of Llama-2-7B and GPT-2-XL.
@@ Appendix D Additional Heatmaps for Pairwise Accuracy Differences
Appendix E Additional Ablation Analysis
We show ablation analysis on GPT-2-XL and Llama-2-7B in Figure 11 and 12.
@@ Appendix E Additional Ablation Analysis
Appendix F Full Table with Varying Number of Optimization Examples for DNIP
The full results for all three models across varying optimization set sizes are shown in Table 14.